OpenStudy (zyberg):

Why 2x^2 + \sqrt{2x^2 + 1} = 11 roots are only +- 2? When I solve I get two more.

OpenStudy (zyberg):

\(2x^2 + \sqrt{2x^2 + 1} = 11\) \(2x^2 = n\) \(n + \sqrt{n + 1} = 11\) \((11 - n)^2 = (\sqrt{n+1})^2\) \(n^2 - 22n + 121 = n + 1\) \(n%2 - 23n + 120 = 0\) \(D = 7^2\) \(n_{1, 2} = \frac{23 +- 7}{2} = 8 and 15\) So, n is 8 and 15, that would give the answer of x to be: \(x = +- 2, +- \sqrt{7.5}\) Where am I wrong?

OpenStudy (princeharryyy):

try plugging back the last two they won't satisfy the equation.

OpenStudy (princeharryyy):

answer would be 19 for them.

OpenStudy (princeharryyy):

and you are not wrong anywhere, one should check if the roots that came out in the final answers satisfy the equation above.

OpenStudy (zyberg):

Oh! That's interesting! Thank you very much! :)

OpenStudy (princeharryyy):

:)